OpenStudy (anonymous):

Let f and g be linear functions with equations f(x) = m1x + b1 and g(x) = m2x + b2. (a) Is f of g also a linear function? b) if so what is the slope

OpenStudy (anonymous):

\[f(g(x))=f(m_2x+b_2)=m_1(m_2x+b_2)+b_1\]

OpenStudy (dumbcow):

a) f(g) = m1(m2x+b2)x +b1 =m1m2x^2 + m1b2x +b1 quadratic not linear

OpenStudy (anonymous):

\[=m_1m_2x+m_1b_2+b_1\]

OpenStudy (dumbcow):

oops haha nevermind ignore me

OpenStudy (anonymous):

says \[f(g(x))\] not \[f\times g(x)\]

OpenStudy (anonymous):

oh so slope it \[m_1m_2\]